OpenStudy (anonymous):

Area of a triangle Solve for h: A= 1/2bh

OpenStudy (anonymous):

Uhh, what's the base? That's just the formula

OpenStudy (anonymous):

That's all the problem says

OpenStudy (anonymous):

The problem just wants you to "rearrange" the formula to solve for the base instead of the way they gave it to you, solved for Area

OpenStudy (anonymous):

oops, I meant solve for h, height

OpenStudy (anonymous):

A = (1/2) b h 2A = bh 2A/b = h

OpenStudy (anonymous):

Same formula as for Area, but rearranged so that you have "h" all by itself.